OpenStudy (anonymous):

Simplify 49 to the negative 1/2 power

OpenStudy (anonymous):

 thats the problem

OpenStudy (diyadiya):

Use this property: \[\huge x^{-m}= \frac{1}{x^m}\] \[\Large 49^{ \frac{-1}{2}}= \frac{1}{49^{\frac{1}{2}}} \] Now \[\large \text{what is } 49^ \frac{1}{2} ?\]

OpenStudy (anonymous):

do u divide 49 by 1/2?

OpenStudy (anonymous):

Ashley, \[49^{1/2} = \sqrt{49}\] the 1/2 is an Exponent.

OpenStudy (anonymous):

oh! okay i get it now! thankyou

OpenStudy (anonymous):

Well done Diyadiya *votes for best response*

OpenStudy (diyadiya):

Sorry @ashleyb i got disconnected and Thankyou :) @Spacelimbus